Could someone show me how to work this out and find the answer. I am lost.
All marble bags from Mary's Marble Company have 3 orange marbles for every 4 red marbles. If a bag has 18 orange marbles, how many red marbles are there in the bag?
Thank you

Answer by ptaylor(2198)

For every 7 marbles, 3 are orange and 4 are red
If the the bag contains 18 (6*3) orange marbles, it must contain 6*4=24 red marbles
CK
3/7 = 18/42
3/7 = 3/7
Another way:
Let x= number of red marbles, then:
3/4 = 18/x
3x=72
x=24
Hope this helps--ptaylor